OpenStudy (anonymous):

sand is falling into a conical pile so that the radius of that base of pile is always equal to half of its height.if the sand is falling at the rate of 10cm3/sec,how fast is height of the pile increasing when the pile is 5cm deep?(topic from calculus 1 related rates)

OpenStudy (anonymous):

the answer is dh/dt = 8/5pie.

OpenStudy (anonymous):

can someone solve it for me..

OpenStudy (amistre64):

we need some derivatives to play with; what formulas can you think of that will relate to the problem at hand?

OpenStudy (amistre64):

perhaps since a volume is being added at a specified rate, we could work with the formula for the volume of a cone?

OpenStudy (anonymous):

Welcome to Open Study and congrats on asking your first question. :)

OpenStudy (amistre64):

we would gladly solve it WITH you, but not FOR you .... we would rather you participate in the solution process please.

OpenStudy (anonymous):

volume of cone is 3/4(pie)(r)squared(h)

OpenStudy (anonymous):

Yes, as @amistre64 said, we can help you, but we will Not just hand you the answers.

OpenStudy (amistre64):

cool, cause i cant recall those things :) \[V=\frac13pi~r^2h\] that looks about right to me ... http://math.about.com/od/formulas/ss/surfaceareavol_2.htm

OpenStudy (amistre64):

the best thing I find to do is just dive in and take an implicit derivative, i see a product rule for the most part

OpenStudy (amistre64):

\[V'=\frac13pi(2rhr'+r^2h')\] and we want to solve for h' \[\frac{\frac 3{pi}V'-2hr~r'}{r^2}=h'\] it should look better with the given values :)

OpenStudy (anonymous):

ooo

OpenStudy (amistre64):

we are given that r = h/2; therefore the rates related are: r' = 1/2 h' as an after thought lets fill that in and then simplify it \[V'=\frac13pi(2rhr'+r^2h')\] \[V'=\frac13pi(2rh\frac12h'+r^2h')\] \[V'=\frac13pi(rh~h'+r^2h')\] \[V'=h'\frac13pi(rh+r^2)\]

OpenStudy (anonymous):

i'll try with using that

OpenStudy (amistre64):

let us know if you get stuck :) i still have difficulties with these things

OpenStudy (anonymous):

yeah my lecturer also have prob with this questions

OpenStudy (anonymous):

he ask us to solve for him instead =.=

OpenStudy (amistre64):

well, as the lecturer he has that right :) we know the givens now as: V' = 10 h = 5 r = 5/2 h' = unknown \[V'=h'\frac13pi(rh+r^2)\] \[10=h'\frac13pi(5(\frac52)+(\frac52)^2)\] \[\frac{30}{pi}=h'(\frac{25}2+\frac{25}4)\] \[\frac{120}{pi}=h'(50+25)\] \[\frac{120}{75pi}=h'\] \[h'=\frac{8}{5pi}\]

OpenStudy (amistre64):

the key for me is in finding suitable formulas or relationships between what is given, and what we want to find.

OpenStudy (anonymous):

wow thx amistre64

OpenStudy (amistre64):

look over it, and try to figure out what i did, with practice it should become more clearer :) good luck